Suppose that the competitive firms are facing the following cost function and revenue function, C = 14 + q2,revenue function: = 20q -q2. (a) Find profit function (b)Find marginal cost function. (c) Find marginal revenue function (c)Find output (q) that maximize profit (d)find total Profit


1
Expert's answer
2022-03-11T08:46:48-0500


a) Profit function

=TR-TC

"20q-q^2-(14+q^2)= 20q-q^2-14-q^2"

"=20q-2q^2-14"

b) Marginal cost

"=\frac{\delta TC}{\delta q}= 2q"

c) Marginal revenue function

"=\frac{\delta TR}{\delta q}= 20-2q"

d) Output that maximizes profit

MR=MC

2q=20-2q

4q= 20

q= 5

e) Total Profit

"=20q-2q^2-14"

"=20(5)-2(5^2)-14= 100-50-14"

"= 36"